# Automatically Generating UI Code from Screenshot: A Divide-and-Conquer-Based Approach

### v2

@@ 5.1.2. Screenshot Subdivision
We divide the screenshot into smaller segments by alternately and recursively applying horizontal and vertical divisions to mirror real-world development practices (Chopper, 2024). Specifically, for each input screenshot, DCGen first attempts to divide the screenshot along horizontal separation lines. If no such lines are detected, it then proceeds to segment the screenshot using vertical separation lines by calling the separation line detector again. This process is applied recursively to each resulting segment until no further separation lines can be identified, or until a user-defined maximum recursion depth is reached. Figure 7 illustrates this subdivision procedure.
@@ 5.2. Assembly
In the assembly process, we integrate the solutions of image subdivisions to generate complete UI code for the original screenshot. The assembly process is essentially the reverse of the division procedure, where code from smaller, child segments is progressively integrated to build up their parent segments. This recursive assembly continues until the full structure of the website is restored.
As shown in Figure 4, for each leaf image segment, DCGen provides the MLLM (i.e., Leaf-solver MLLM) a website screenshot with a red rectangular bounding box specifying the part of the image it should focus on, then ask it to describe the image segment. For each parent image segment, DCGen provides the MLLM (i.e., Assembly MLLM) a screenshot with a bounding box, as well as all the descriptions generated from its children’s image segments. In the last step (merging for root node), DCGen calls the Generation MLLM to produce the complete UI code for full screen, based on the descriptions from its children segments and the entire screenshot.
We adapt the basic instructions based on the previous study by Si et al. (Si et al., 2024). All the prompt details for the leaf-solver MLLM, assembly MLLM, and generation MLLM are made available in our artifact.
@@ 6. Experiment
  • RQ1: How does DCGen compare to various prompting strategies in design-to-code?
  • RQ2: How robust is DCGen for complex designs?
  • RQ3: How generalizable is DCGen for other state-of-the-art (SOTA) MLLMs?
  • RQ4: What is human’s feedback on DCGen in real-world application development workflows?
@@ 6.1.1. Backbone Models
To assess the performance of DCGen, we employ it to evaluate three SOTA MLLMs: GPT-4o (OpenAI, 2024a), Gemini-1.5 (Google, 2024), Claude-3 (Anthropic, 2024a). Access to all models is facilitated through their respective official APIs, the specific model numbers are 20240513 for GPT-4o, 20240229 for Claude-3-Sonnet, and Gemini-1.5-flash-latest accessed during May 2024. Due to a limitation on computational resources, we randomly sampled 120 websites from our dataset to form the testing set. Table 3 shows the statistics of the testing dataset, the data is rich in HTML tags and diverse in overall complexity.
@@ 6.1.2. Parameter Setting
For MLLM model configurations, we set the temperature to 0, and the $max\_tokens$ parameter to the maximum allowable value for each model (i.e., 4096 for GPT-4o and Claude3-Sonnet, 8192 for Gemini-1.5-Flash), while other parameters are maintained at their default settings as specified in the corresponding API documentation (Google AI, 2024; OpenAI, 2024b; Anthropic, 2024b). For parameters in the image segmentation algorithm detailed in Section 5.1, we set var_thr=60, diff_thresh=5, diff_portion=0.3, and set the division depth to one.
@@ 6.1.3. Efficiency
Our experiments were conducted on a Windows 10 PC with an IntelCore i7-10870H CPU. We assess the performance of DCGen by first evaluating the speed of the Image subdivision algorithm, which completes the subdivision of a 1920x1080 pixel image in less than one second when the division depth is set to one. On average, DCGen generates code for an entire webpage (1,370 tags) within 4 minutes. This demonstrates DCGen’s practical efficiency in accelerating web development.
@@ 6.1.4. Baselines
Chain-of-Thought (CoT) prompting (Wei et al., 2022a) generates a chain of thought for each question and then generates the corresponding code. For CoT, we use the ”let’s think step by step” instruction from Chae et. al (Chae et al., 2024). Self-refine prompting (Chen et al., 2023) let the model refine its own generated code via multi-turn conversation. We adopt the self-refine prompting and direct promoting method from Si et al. (Si et al., 2024). We do not select previous deep-learning-based approaches as baselines since they are either tailored to the mobile app environment (Chen et al., 2018; Nguyen and Csallner, 2015; Moran et al., 2018a), not opensourced (Xu et al., 2021), do not take website screenshot as input or do not produce complete website code (Asiroglu et al., 2019; Beltramelli, 2018). The details of prompting strategies used in our experiment as baselines are all available in our artifact.
Table 3. Statistics of the test data
Min Max Average
Length (tokens) 338 992,008 41,264
Tag Count 1 397 175
DOM Depth 1 34 12
Unique Tags 1 57 17
Total size 120
@@ 6.2. Automatic Metrics
Let $C_{0}$ be the original HTML+CSS code of the website, $I_{0}$ be the screenshot of the original website, and $C_{g}$, $I_{g}$ be the code and screenshot of the generated website. We evaluate the performance of DCGen in both high-level similarity and fine-grained element matching.
@@ 6.2.1. High-Level Similarity
For code similarity, we adopt the Bilingual Evaluation Understudy (BLEU) score (Papineni et al., 2002) to evaluate the closeness of generated code $C_{g}$ to the original code $C_{0}$. BLEU score is an automatic evaluation metric widely used in machine translation studies. It calculates the similarity of machine-generated translations and human-created references (i.e., ground truth). It is also widely used in code generation tasks to measure the similarity between generated code and the ground truth  (Chen et al., 2021; Yin et al., 2018; Wei et al., 2019; Ren et al., 2020; Lu et al., 2021).
For image similarity, we measure image similarity between $I_{0}$ and $I_{g}$ via CLIP Score  (Si et al., 2024). CLIP score relies on the similarity of their CLIP (Radford et al., 2021) embeddings, denoted as CLIP($I_{0}$, $I_{g}$). Specifically, we extract features using CLIP-ViT-B/32 after resizing the screenshots to squares.
@@ 7. Results & Analysis
7.1. RQ1: How does DCGen compare to various prompting strategies?
@@ 7.1.1. High-level Performance
Figure 9 presents the overall performance of various methods evaluated on the CLIP and BLEU metrics tested on GPT-4o. The results indicate that DCGen achieves the most substantial performance gains, outperforming the direct prompting for 13.42% and 1.39% for CLIP and BLEU scores, respectively. Such promising performance demonstrates the effectiveness of our divide-and-conquer approach. Besides, while DCGen and the self-refine approach contribute positively to model performance, the Chain-of-Thought (CoT) prompting approach has a negative impact on the model’s capabilities. A detailed analysis of the code generated under CoT instructions reveals that 70% (84 out of 120) of the HTML files contain no CSS, leading to purely textual websites without color or style. Further investigation suggests that the model, when focused on step-by-step visual reasoning instructions, neglected the directive to include both HTML and CSS in a single output, indicating a limitation in handling large amounts of information simultaneously.
@@ 7.1.2. Fine-grained Metrics
[…]
@@
Answer to RQ1: DCGen consistently outperforms other design-to-code techniques on both high-level and fine-grained metrics, due to the precise and focused image segments.
Method CLIP BLEU
Direct 69.94 6.78
CoT 62.52 6.20
Self-Refine 75.19 7.91
DCGen (Ours) 83.36 8.17
Figure 8. DCGen’s high-level performance on GPT-4o (%).
(a) Fine-grained metrics.
(b) Error rate.
Figure 9. Detailed analysis.
7.2. RQ2: How robust is DCGen for complex designs?
Recognizing the diverse complexity of real-world websites, this RQ evaluates DCGen’s efficacy on websites with varying complexity levels. We quantify website complexity using the number of HTML tags and focus on changes in visual metrics (i.e., CLIP score and the average of all four fine-grained metrics) using GPT-4o as the backbone. In our sample dataset, the number of HTML tags varies from 24 to 298, where a higher number of tags indicates increased complexity in the websites.
Figure 10(a), and Figure 10(b) showcase DCGen’s performance across varying degrees of complexity. DCGen consistently achieves leading performance across all metrics, demonstrating its strong ability to maintain high visual scores over varying complexity levels. Specifically, as website complexity escalates from 24 tags to 298 tags, DCGen’s performance remains stable, with only a minor 5.1% decline in effectiveness as measured by the CLIP score. Even though the division stage is straightforward, such consistent performance underlines the robustness and effectiveness of DCGen in handling website complexity.
Answer to RQ2: DCGen delivers superior performance across various website complexities, indicating its robustness in handling both simple and intricate webpages.
(a) CLIP score vs. website complexity.
(b) Avg. fine-grain score vs. website complexity.
Figure 10. DCGen’s robustness for complex websites.
7.3. RQ3: How generalizable is DCGen for other SOTA MLLMs?

### v3

@@ 5.1.2. Screenshot Subdivision
We divide the screenshot into smaller segments by alternately and recursively applying horizontal and vertical divisions to mirror real-world development practices (Chopper, 2024). This process is applied recursively to each resulting segment until no further separation lines can be identified or until a user-defined maximum recursion depth is reached.
@@ 5.2. Assembly
In the assembly process, we integrate the solutions of image subdivisions to generate complete UI code for the original screenshot. The assembly process is essentially the reverse of the division procedure, where code from smaller child segments is progressively integrated to build up their parent segments until the full structure of the website is restored.
We propose two different implementations of the assembly process, agent-based and rule-based. Agent-based assembly is shown in Figure 3. DCGen first provides the MLLM (i.e., Leaf-solver MLLM) with leaf image segments, then asks it to generate the code of the segments. Then, DCGen provides the MLLM (i.e., Assembly MLLM) a parent segment and all the code generated from its children’s image segments. In the last step (merging for root node), DCGen calls the Generation MLLM to produce the complete UI code for full screen based on the code from its children segments and the entire screenshot. We adapt the basic instructions based on the previous study by Si et al. (Si et al., 2024). All the prompt details for the leaf-solver MLLM, assembly MLLM, and generation MLLM are made available in our artifact. Algorithm 2 shows the pseudocode of the complete DCGen algorithm.
For rule-based assembly, instead of using multiple MLLMs to assemble partial code output at different depths, we replace the assembly step with a rule-based program that uses CSS grid for layout positioning [^4] https://www.w3schools.com/css/css_grid.asp. As usual, we first divide the input image into multiple segments. The position of each segment within the overall layout is recorded. Then, MLLMs are used only at the leaf level to independently generate HTML and CSS code for each segment. Consequently, instead of using LLMs to assemble the segments into a full webpage, we directly place each generated segment into a CSS grid layout at the recorded positions. The assembled template, now containing all the fine-grained code blocks in the correct positions, is passed to a generator LLM for final adjustments, ensuring cohesion, flexibility, and accurate visual details. Algorithm 3 shows the pseudocode of this implementation.
Algorithm 2 DCGen-Agent
1:Input: $img$, $bot$, $depth$, $max\_depth$, $leaf\_prompt$, $node\_prompt$
2:if $depth\geq max\_depth$ then
3:    return  $bot.generate\_code(leaf\_prompt,img)$
4:end if
5:$img\_cuts\leftarrow segment\_image(img)$
6:if no cuts found then
7:    return  $bot.generate\_code(leaf\_prompt,img)$
8:end if
9:$code\_parts\leftarrow[]$
10:for each $cut$ in $img\_cuts$ do
11:    $code\_parts\leftarrow$ run in new thread: $dcgen(cut,bot,depth+1,max\_depth,leaf\_prompt,node\_prompt)$
12:end for
13:
14:return  $bot.generate\_code(node\_prompt+code\_parts,img)$
Algorithm 3 DCGen-Rule
1:Input: $img$, $bot$
2:$img\_segs,seg\_positions\leftarrow segment\_image(img)$
3:$code\_segs\leftarrow generate\_code\_parallel(img\_segs,bot)$
4:$HTML\_template\leftarrow CSS\_grid(seg\_positions,codes\_segs)$
5:$final\_code\leftarrow generate\_code(HTML\_template,img,bot)$
6:
7:return  $final\_code$
Figure 5. Statistics of the test data
Min Max Average
Length (tokens) 383 123,578 28,430
Tag Count 2 396 188
DOM Depth 2 28 12
Unique Tags 2 57 19
Total size 111
@@ 6. Experiment
  • RQ1: How do different parameters influence the performance of the separation algorithm?
  • RQ2: How does DCGen compare to various methods in design-to-code?
  • RQ3: How robust is DCGen for complex designs?
  • RQ4: How generalizable is DCGen for other state-of-the-art (SOTA) MLLMs?
  • RQ5: What is human’s feedback on DCGen in real-world application development workflows?
@@ 6.1.1. Backbone Models
To assess the performance of DCGen, we employ it to evaluate four SOTA MLLMs: GPT-4o (OpenAI, 2024a), Gemini-1.5-flash and Geimini-2-flash (Google, 2024), Claude-3-Sonnet (Anthropic, 2024a). Access to all models is facilitated through their respective official APIs. The specific model numbers are 20240513 for GPT-4o, 20240229 for Claude-3-Sonnet, Gemini-1.5-flash-latest accessed during May 2024, and Gemini-2.0-flash accessed Feb 2025. Due to a limitation on computational resources and budget, we randomly sampled 111 websites from our dataset to form the testing set. Table 5 shows the statistics of the testing dataset. The data is rich in HTML tags and diverse in overall complexity.
@@ 6.1.2. Parameter Setting
For MLLM model configurations, we set the temperature to 0 and the $max\_tokens$ parameter to 8192 for each model, while other parameters are maintained at their default settings as specified in the corresponding API documentation (Google AI, 2024; OpenAI, 2024b; Anthropic, 2024b). For parameters in the image segmentation algorithm detailed in Section 5.1, we set var_thr=50, diff_thresh=45, diff_portion=0.9, window_size=50 and set the division depth to 2.
@@ 6.1.3. Efficiency
DCGen leverages parallelism to promote efficiency, where all queries at the same depth are processed simultaneously, as shown in Algorithm 2. Given an MLLM response time of $t$ per query, the total response time for DCGen is $(depth+1)*t=3t$ for $depth=2$. As for cost efficiency, on each depth where MLLM produces partial code for an image segmentation, the sum of the output code on each depth is roughly upper bound by the length of the final generated code $L$. Thus, the total token usage (and cost) for agent-based assembly is upper bounded by $(depth+1)*L=3L$ for $depth=2$. In short, DCGen costs three times that of direct querying. However, with the rule-based assembly, DCGen is able to eliminate intermediate MLLM queries by generating code at the leaf and root levels. In such a way, the optimal cost can be lowered to $2L$, which stays the same as self-refinement prompting.
@@ 6.1.4. Baselines
We first compare our method with the Pix2Code method (Beltramelli, 2018) that trained deep-learning model to convert screenshot to code. Then we compare various prompting strategies including Chain-of-Thought (CoT) prompting (Wei et al., 2022a) that generates a chain of thought for each question and then generates the corresponding code. For CoT, we use the ”let’s think step by step” instruction from Chae et. al (Chae et al., 2024). Self-refine prompting (Chen et al., 2023) let the model refine its own generated code via multi-turn conversation. We adopt the self-refine prompting and direct promoting method from Si et al. (Si et al., 2024). The details of prompting strategies used in our experiment as baselines are all available in our artifact.
@@ 6.2. Automatic Metrics
Our experiments use screenshots of existing webpages as proxies for design images $I_{0}$, and use their corresponding HTML+CSS code as ground truth $C_{0}$. We evaluate the performance of DCGen in both high-level similarity and fine-grained element matching.
@@ 6.2.1. High-Level Similarity
For code similarity, we adopt the Normalized Levenshtein Distance (edit distance) (RapidFuzz, 2024) to evaluate the closeness of generated code $C_{g}$ to the original code $C_{0}$. Levenshtein Distance is a robust method that is widely used to measure similarity between two sequences (Haldar and Mukhopadhyay, 2011). The score is calculated as S = $1-distance/(l_{1}+l_{2})$ (RapidFuzz, 2024), which provides a direct measurement of similarity between the generated and ground truth HTML, with a higher value indicating higher similarity.
For image similarity, we measure image similarity between $I_{0}$ and $I_{g}$ via CLIP Score  (Si et al., 2024). CLIP score relies on the similarity of their CLIP (Radford et al., 2021) embeddings, denoted as CLIP($I_{0}$, $I_{g}$). Specifically, we extract features using CLIP-ViT-B/32.
@@ 7. Results & Analysis
7.1. RQ1. How do parameters influence the performance of the separation algorithm?
In this section, we examine how five hyper-parameters, i.e., diff_thresh, var_thresh, diff_portion, window_size, and depth, affect the performance of the separation algorithm. To evaluate the precision of the separation algorithm, we extract bounding boxes formed by the separation lines and use the standard Intersection over Union (IoU) metric (Redmon et al., 2015) to quantifies the overlap between predicted and ground-truth bounding boxes. The IoU between two bounding boxes $b_{1},b_{2}$ is calculated as $IoU=\frac{\text{Intersection}(b_{1},b_{2})}{\text{Union}(b1,b2)}$, with higher value indicating more overlapping or higher precision.
For each parameter configuration, we assess the algorithm’s performance using our entire test set. Specifically, given an input image, we recursively apply our separation algorithm to generate separation lines. The rectangular areas defined by these lines serve as predicted bounding boxes. We then extract ground-truth bounding boxes from the corresponding HTML code using JavaScript. For each predicted bounding box, we compute its maximum IoU with all ground-truth bounding boxes. The average IoU across the test set provides a measure of the algorithm’s precision. However, it is important to note that IoU serves only as a reference metric. The goal of the separation algorithm is not to perfectly match each HTML element but to generate reasonable divisions that help MLLMs focus on finer details.
Beyond IoU, we also evaluate the extent of image division. Some parameter settings may yield highly accurate but overly coarse segmentations (e.g., producing only a few bounding boxes), while others may result in more granular yet less precise divisions. To quantify this, we introduce the Separation $\text{Separation Rate}=1-\frac{b_{max}}{S}$, where $b_{max}$ represents the size of the largest undivided leaf bounding box, and $S$ is the total image size. A higher separation rate indicates finer image segmentation.
7.1.1. Effect of diff_thresh, var_thresh, and diff_portion.
To analyze the impact of diff_thresh, var_thresh, and diff_portion, we fixed depth=2 and window_size=50, then selected three diff_portion values (0.3, 0.6, and 0.9) spanning its range (0,1). For each diff_portion, we plotted the IoU scores across different diff_thresh and var_thresh settings.
Figure 6(a) shows that diff_portion and diff_thresh have minimal impact ($\leq 1\%$) on algorithm precision, whereas a lower var_thresh improves accuracy. Experiments indicate that IoU variations across parameter settings remain within approximately 5%, demonstrating the separation algorithm’s robustness. These fluctuations arise because lower var_thresh reduces the likelihood of misclassifying areas as “blank”, thereby limiting unnecessary separation proposals, which results in a more conservative approach, yielding fewer but more accurate separation lines. Meanwhile, as illustrated in Figure 6(b), increasing var_thresh significantly enhances the separation rate, producing finer-grained segmentations.
Figure 6. (a) Effect of diff_portion, diff_thresh, and var_thresh on the separation algorithm’s IoU score. (b) Effect of diff_thresh, var_thresh on separation rate when diff_portion=0.9. Larger var_thresh improves separation rate.
7.1.2. Influence of Window Size and Depth.
To analyze the effect of sliding window size and separation depth on the separation algorithm, we selected three sets of parameters from Section 7.1.1 and plotted the average IoU across different window sizes and depths. The results, shown in Figure 7(a), indicate that a smaller depth and a larger window size generally lead to more accurate separation.
On the other hand, we observe a similar trade-off between precision and separation rate. Figure 7(b) shows that a larger depth and smaller window size lead to a finer separation rate. However, while an increase from 1 to 2 enhances the separation rate, further increases in depth barely influence the rate.
Figure 7. (a) IoU under different window sizes and depths across three sets of var_thresh (vt), diff_thresh (dt), and diff_portion (dp). (b) Separation rate under different window sizes and depths when vt=5, dt=50, and dp=0.9. A larger depth and smaller window size lead to a finer separation rate. Depth 2, 3, and 4 have a similar separation ratio.
7.1.3. Parameter Selection
We selected the parameters based on observations from Sections 7.1.1 and 7.1.2. Observing that diff_portion and diff_thresh have minimal influence on algorithm precision, we randomly pick diff_portion=0.9 and diff_thresh=45 for implementing DCGen. While smaller var_thresh improves separation accuracy, it fluctuates by only about 5%. On the other hand, larger var_thresh significantly improves the separation rate. Therefore, we pick a larger var_thresh=50. Then, to achieve a better separation rate and maintain a reasonable IoU, we pick window size=50. Finally, since an increase of depth from 1 to 2 obviously enhances the separation rate, further increases in depth barely influence the rate but decrease IoU significantly, we pick depth=2 to achieve a balance between precision and division rate. Figure 8 shows the trade-off we observed in Sections 7.1.1 and 7.1.2, and Figure 8(c) shows an example division of the parameters we adopted.
Answer to RQ1: Diff portion and diff thresh have minimal influence on the algorithm; Smaller vt, shallower depth, and larger ws produce fewer but larger and more accurately segmented areas; Larger vt, greater depth, and smaller ws generate smaller and finer divisions, albeit with reduced accuracy.
7.2. RQ2: How does DCGen compare to various methods?
@@ 7.2.1. High-level Performance
Table 5 presents the high-level performance comparison of DCGen and several baseline methods across the CLIP and Code metrics. DCGen outperforms all other methods, achieving a notable 85.62% on CLIP and 15.34% on Code. Specifically, it surpasses the self-refine approach by 1.66% on CLIP and 1.69% on Code, highlighting its superior design-to-code generation capabilities. Notably, the Chain-of-Thought (CoT) method, while improving the text generation (Code: 13.64%), does not significantly outperform other methods in visual fidelity.
@@ 7.2.2. Fine-grained Metrics
Table 5 showcases the fine-grained performance comparison of the methods in terms of Block-Match, Text, Position, and Color metrics. DCGen excels in these areas, achieving the highest scores for Block-Match (81.7%), Text (97.8%), Position (83.6%), and Color (85.4%). The results suggest that DCGen provides a more accurate and complete design-to-code transformation, with particularly strong improvements in color accuracy (3.5%) and text generation (1.6%).
(a) Input image
(b) IoU=0.58, Sep Rate=0.45
(c) IoU=0.43, Sep rate=0.61
Figure 8. Comparison of different parameter settings for the separation algorithm. (b) ws=90, vt=1, dt=45, dp=0.9, depth=1. A larger ws, smaller vt and depth result in fewer but more accurate segmentation areas (c) ws=50, vt=50, dt=45, dp=0.9, depth=2. A smaller ws, larger vt and depth lead to finer but less precise segmentation (these are the parameters we used in the final experiment).
@@
Answer to RQ2: DCGen consistently outperforms other design-to-code techniques on both high-level and fine-grained metrics, due to the focused image segments.
Table 3. DCGen’s performance on GPT-4o (%).
Method CLIP Code
Direct 83.72 12.76
CoT 83.64 13.64
Self Ref. 83.96 13.65
DCGen 85.62 15.34
Pix2Code 73.35 9.70
Table 4. Fine-grained comparison of different methods (%).
Method Block-Match Text Position Color
Direct 76.1 95.9 81.6 81.8
CoT 79.5 88.4 76.9 76.2
Self Ref. 81.4 96.1 82.7 81.9
DCGen 81.7 97.8 83.6 85.4
Table 5. DCGen + GPT-4o on Design2Code-Hard (%).
Method CLIP Code
Direct 81.04 14.37
CoT 79.97 14.90
Self Ref. 80.43 15.20
DCGen 84.02 19.48
Pix2Code 73.56 11.17
7.3. RQ3: How robust is DCGen for complex designs?
To evaluate the limits of DCGen, we test it on the Design-to-Code-Hard dataset, which consists of 80 highly complex webpages from GitHub Pages. This dataset presents significant challenges for state-of-the-art MLLMs in converting visual designs into code (Si et al., 2024). Despite the increased difficulty, DCGen achieves the best performance in both code and visual similarity, as shown in Table 5.
To further understand DCGen’s role in improving results on complex webpages, we quantify website complexity using multiple metrics: image size, DOM depth, and the number of HTML tags, and focus on visual metrics (CLIP score) and code-generation metrics. We further include the Design-to-Code-Hard dataset, comprising 80 particularly challenging webpages, alongside our original dataset to provide a comprehensive assessment. The results measured by DOM depth and the number of HTML tags are presented on our website, which shows a moderate improvement of around 2% across different complexities. The results measured by image size are summarized in Table 6:
  1. (1)
    The separation rate remains high ($\geq$50%) across various complexity levels.
  2. (2)
    IoU scores are relatively stable (30%-40%) across most complexity scenarios, with a notable dip in the image size interval (15, 20].
  3. (3)
    When measuring complexity by image size, DCGen consistently outperforms existing prompting methods, particularly for large images (15M-25M pixels), despite the IoU dip observed.
These findings confirm that DCGen primarily benefits the conversion of large design images into code by breaking complex screenshots into smaller, more manageable visual segments. Even with imperfect segmentation, this strategy helps MLLMs better concentrate on detailed sub-tasks, significantly reducing omissions, distortions, and misarrangements. This finding is consistent with our motivating study, which revealed that decomposing large screenshots into smaller visual components improves model performance by facilitating focused reasoning steps.
Answer to RQ3: DCGen is robust to complex websites. It primarily benefits the conversion of large design images into code by breaking complex screenshots into smaller, more manageable visual segments.
Table 6. Performance comparison across different input image sizes (in million pixels). All metrics are scaled by 100. Promptmax denotes the maximum score among self-refine, CoT, and direct prompting. “DCGen Advantage” indicates performance improvement of DCGen over Promptmax, with positive values highlighted in green and advantages greater than 5 underlined.
Image size (M) (0, 5] (5, 10] (10, 15] (15, 20] (20, 25]
Sep Rate 58.8 70.8 75.0 84.1 51.4
IoU 43.6 41.5 36.5 21.2 45.1
CLIP Promptmax 84.7 83.4 76.1 72.9 75.2
CLIP DCGen 86.3 83.9 80.2 81.3 82.7
DCGen Advantage 1.6 0.5 4.1 8.4 7.5
Code Promptmax 15.9 11.3 5.5 11.5 4.5
Code DCGen 18.2 15.7 5.9 27.1 6.5
DCGen Advantage 2.3 4.4 0.4 15.6 2.0
7.4. RQ4: How generalizable is DCGen for other SOTA MLLMs?